# SABR-Informed Multitask Gaussian Process: A Synthetic-to-Real Framework for Implied Volatility Surface Construction

### v1

@@ 5.1 Performance in the Base Scenario
Figure 3: Implied volatility curve comparison and residuals under the Base Heston scenario for (a) Near-term, (b) Mid-term, and (c) Long-term maturities.
The results showed different performance patterns across the term structure. For near-term maturity (Figure 3(a)), where market data is relatively dense, SABR-MTGP achieved the most accurate fit across strikes (RMSE=0.0014). The standard GP (RMSE=0.0024) showed systematic deviations, suggesting difficulty capturing the precise smile pattern. Direct SABR interpolation (RMSE=0.0057) struggled with curvature, showing systematic errors consistent with limitations of its parametric form in matching these Heston dynamics.
Moving to mid-term maturity (Figure 3(b)), data sparsity increased. SABR-MTGP continued to provide the closest fit (RMSE=0.0005). SABR interpolation also performed very well (RMSE=0.0009), capturing the overall shape effectively. However, the standard GP (RMSE=0.0043) showed larger and more scattered errors, particularly for strikes away from the money.
For long-term maturity (Figure 3(c)), SABR interpolation performed best according to RMSE (0.0005), benefiting from its structure to provide a reasonable approximation. SABR-MTGP also achieved very good accuracy (RMSE=0.0007). The standard GP (RMSE=0.0025) exhibited notably higher errors compared to the other two methods, with residuals spanning approximately from $-0.006$ to $+0.002$ volatility points, indicating difficulties in generalizing from the sparse long-term data.
These findings highlight the advantages of our hybrid approach. In data-rich regions (near-term), SABR-MTGP combined market data with structural guidance to outperform alternatives. As data gets sparser (mid- to long-term), it used SABR structure effectively while retaining flexibility to adapt to Heston dynamics, balancing SABR’s rigidity and standard GP’s data dependence.
5.2 Robustness Across Market Conditions
The performance heatmaps in Figures 4, 5, and 6 summarized model accuracy (RMSE and MeanAE) across the ten Heston parameter settings, evaluated at near-, mid-, and long-term horizons. Lighter shades indicated lower errors.
Figure 4: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Near-term maturity ($\tau=0.3$).
Figure 5: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Mid-term maturity ($\tau=0.9$).
Figure 6: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Long-term maturity ($\tau=2.2$).
These heatmaps show consistent performance patterns. SABR-MTGP consistently demonstrated strong performance, frequently ranking as the best or a very close second-best method. The standard GP showed reasonable performance in the near-term, but its accuracy dropped significantly at long-term maturities, where it yielded the highest errors. Conversely, SABR interpolation struggled in the near-term, often producing the highest errors, but benefited from its fixed structure in long-term cases. It frequently achieved the lowest errors, especially with very sparse data. SABR-MTGP’s ability to adaptively blend structural guidance with data-driven flexibility allowed it to maintain high accuracy overall, overcoming SABR’s near-term issues and the GP’s long-term instability.
@@ 5.3 Learned Task Relationship
The MTGP framework learned the relationship between the SABR source task and the Heston market target task. Figure 7 shows this adaptive relationship for three Heston settings, illustrating variations in learned correlation, covariance structure, and variance decomposition (task-specific proportion $\kappa^{2}_{\mathcal{Z}}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$ vs. shared variance proportion $\sigma^{2}_{h}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$).
(a) Low Correlation Example: Moderate Mean-Rev Setting (Learned Correlation = 0.22).
(b) Moderate Correlation Example: Base Setting (Learned Correlation = 0.41).
(c) High Correlation Example: Term Structure Up Setting (Learned Correlation = 0.90).
Figure 7: Examples of learned task relationship analysis for the SABR-MTGP model under different Heston parameter settings: (a) Moderate Mean-Rev, (b) Base, and (c) Term Structure Up. Panels show variations in learned correlation (left), covariance structure (center), and variance decomposition (right).
The learned correlation varied with Heston parameters. Faster mean reversion (Moderate Mean-Rev) yielded low correlation (0.22), perhaps because SABR captured these dynamics less well. An upward term structure (Term Structure Up) resulted in high correlation (0.90), suggesting SABR provided a better structural approximation in this case. The Base setting showed moderate correlation (0.41). This showed the model’s ability to assess similarity between SABR information and target data based on market conditions. The task covariance matrices reflected these correlations.
A key pattern was that the shared covariance component consistently accounted for most variance in both tasks across scenarios. Even if SABR is not a perfect match (low correlation), the MTGP leveraged shared properties like smoothness, preventing overfitting to sparse target data and performing better than the standard GP at longer maturities. When SABR aligns well (high correlation), the model strongly used this information, achieving accuracy competitive with or better than SABR interpolation, while maintaining flexibility.
5.4 Parameter Sensitivity Analysis
We further examined robustness by analyzing performance sensitivity to Heston parameter changes. Using the coefficient of variation (CV) of RMSE, Figure 8 showed how much the precision of each model fluctuated within the parameter categories relative to its average performance. Lower CV (lighter colors) indicated more stable performance.
(a) Near-term ($\tau=0.3$).
(b) Mid-term ($\tau=0.9$).
(c) Long-term ($\tau=2.2$).
Figure 8: Parameter sensitivity heatmaps showing the Coefficient of Variation (CV) of RMSE for GP, SABR, and SABR-MTGP across different Heston parameter categories (y-axis) for near-, mid-, and long-term evaluation maturities. Lower CV (lighter color) indicated greater performance stability within that parameter category.
The sensitivity heatmaps highlighted model stability differences. The standard GP often showed high CV, indicating its performance was sensitive to underlying volatility dynamics, especially concerning ‘Correlation’ and ‘Vol-of-Vol’ changes at longer maturities. SABR interpolation typically had the lowest CV due to its fixed parametric form, but this stability could come at the cost of lower accuracy, particularly near-term. SABR-MTGP generally displayed intermediate stability. It demonstrated improved stability over the standard GP in the mid- and long-term. Although it was less stable than SABR interpolation. The sensitivity results thus complemented accuracy findings: SABR-MTGP found a middle ground, using the prior for improved stability over GP, while retaining adaptive flexibility unlike SABR interpolation.
@@ 5.5 Summary
The experiments using Heston data suggested SABR-MTGP effectively constructed the IVS. By combining SABR’s structural information with market observations in a multi-task framework, it achieved better accuracy and stability than using either a standard GP or SABR. It effectively integrated structural information with observed data, adaptively leveraging SABR insights by assessing their relevance to the target dataset.
@@ 6 Application to SPX Market Data
To further evaluate the practical applicability of SABR-MTGP, we tested its performance on real-world market data for SPX options (provided by the OptionMetrics database), observed on August 1, 2023. This case study complements the controlled experiments using the Heston model by assessing the method’s behavior under actual market conditions.
@@ 6.1 Data
The dataset included European call options on the SPX index. We preprocessed the data by filtering for valid quotes, calculating time-to-maturity, and finding implied volatilities from mid-prices. We used the Zero Coupon Bond yield curve as the interest rate curve, and estimated the dividend yield using put-call parity. This gave us several hundred valid market implied volatility observations across different strikes and maturities up to about one year. Key statistics summarizing this dataset are in Table 2. This processed market data was the target task ($\mathcal{D}_{\mathcal{T}}$) for both the SABR-MTGP model.
@@ 6.2 Results
We do not have a ground truth IVS for real market data. So, the evaluation used a qualitative assessment of the model fits and the resulting surfaces. Figure 9 shows typical implied volatility slices for near-term ($\tau=0.2190$ years) and mid-term ($\tau=0.8898$ years) maturities.
@@
Figure 9: Implied volatility slices for SPX options on August 1, 2023. Comparison of SABR-MTGP, GP, and SABR interpolation against market data points. Residuals (Model - Market) are shown below each slice.
In both slices (Figure 9), all three methods generally captured the observed smile/skew pattern of the market. But the residual plots showed important differences in fit quality. For the near-term slice (Figure 9(a)), SABR interpolation had systematic errors. It underestimated volatility for low strikes (ITM) and overestimated near the money. This was shown by the consistent pattern in its residuals. Both the standard GP and SABR-MTGP fitted the observed market points better, especially around the at-the-money region. But there was a notable difference in the deep in-the-money region (strikes 3250-4000). The standard GP curve, although sloping downward, showed a shape considered less regular and potentially less financially realistic compared to volatility curve given by the SABR. The SABR-MTGP curve maintained a smoother, decreasing volatility in this data-sparse region. This produced a more financially realistic interpolation, likely because of the structural guidance from the SABR source task. For the mid-term slice (Figure 9(b)), where the smile was flatter and the market data was denser, all models achieved a very good fit. Notably, the SABR model fit improved significantly compared to the near-term slice. The residuals for both GP and SABR-MTGP were particularly small (generally within +/- 0.001), demonstrating excellent agreement with market observations.
@@
Figure 10: SPX Implied Volatility Surfaces constructed using (a) GP, (b) SABR Interpolation, and (c) SABR-MTGP, based on market data from August 1, 2023. Black dots represent observed market implied volatilities.
The three plots (Figure 10) showed the different shapes of the implied volatility surfaces produced by the three models. The standard GP surface (Figure 10(a)) matched the observed market data points very closely. However, in data-sparse regions, particularly for deep out-the-money options ($K$ is large), it showed some unusual patterns that were inconsistent with typical market behavior, such as a downward slope in volatility where an upward sloping skew is generally expected. The SABR surface (Figure 10(b)) was naturally smooth because it was based on a specific mathematical formula. But, because of this fixed structure, it might not have captured all the specific details or patterns seen in the market data. The SABR-MTGP surface (Figure 10(c)) seemed to offer a good balance between the other two methods. It fitted the market data well, similar to the standard GP. And, it maintained a smoother and more economically consistent shape compared to the standard GP, especially when extending to areas with less data, such as longer times-to-maturity or strike prices far from the current price. This visual comparison suggested that the SABR-MTGP model successfully used the general structure of the SABR model to guide the construction of the surface, particularly where market data was sparse. This helped to create a more stable and realistic surface overall.
In summary, the application to real SPX market data showed the model’s ability to integrate structural financial knowledge with sparse empirical observations. This led to well-behaved and plausible implied volatility surfaces suitable for practical applications.
@@ 7 Conclusion
In this paper, we proposed SABR-MTGP, an approach bridging structural and data-driven methods for IVS construction. The key idea is framing IVS construction as a multi-task learning problem where we used structural information from the SABR model to improve predictions from sparse market data.
Our approach offered several contributions. First, we show how to combine financial models with machine learning techniques. We used SABR as a complementary information source within our multi-task framework. Second, we developed a hierarchical Bayesian regularization for task embeddings that adaptively determined the level of information transfer. Third, our evaluation using Heston ground truth data showed this approach balanced structure and flexibility. It provided reliable predictions across market conditions and outperformed standard Gaussian process particularly in data-sparse regions. Improvements were notable for longer maturities and extreme strikes, where market data was limited. Finally, an application to real SPX market data further illustrated the model’s ability to produce smooth, reliable IVS that respect market observations while benefiting from structural guidance.
Our work has practical implications for quantitative finance. SABR-MTGP provides a more reliable tool for IVS construction, crucial for derivative pricing, hedging, and risk management. By incorporating structural knowledge and empirical observations, it reduced model misspecification risk while maintaining adaptability to market realities. Future research could extend the multi-task framework to other structural models or multiple models simultaneously.

### v2

@@ 5.1 Performance and Robustness Across Market Conditions
Figure 3: Implied volatility curve comparison and residuals under the Base Heston scenario for near-term
Figure 4: Implied volatility curve comparison and residuals under the Base Heston scenario for long-term
In the long-term slice, the three methods are closer. SABR is often the best while SABR-MTGP remains very close and smooth. The single-task GP shows mild undulations yet the deviations are small.
These patterns align with the design of SABR-MTGP. Short maturities contain more observations and the SABR-based source task acts as a stabilizer that suppresses small oscillations of a pure GP while preserving flexibility. At long maturities data are sparse and a simple parametric structure can be competitive. SABR-MTGP uses learned task embeddings with hierarchical regularization to adapt the level of transfer. When SABR and the target regime align, the learned correlation increases (cf. Section 5.3).
5.2 No-Arbitrage Analysis
We evaluate no-arbitrage by computing the butterfly spread and total variance. For each maturity $\tau$ and ordered strikes $\{K_{i}\}_{i=1}^{100}\subset[80,140]$, we define the butterfly spread
$B_{i}(\tau)\,=\,C(K_{i+1},\tau)\,-\,2\,C(K_{i},\tau)\,+\,C(K_{i-1},\tau),$ (27)
which must satisfy $B_{i}(\tau)\geq 0$ (convexity in $K$). We report the butterfly violation rate as $\{i:B_{i}(\tau)<-\varepsilon\}/\#\{i\}$ with tolerance $\varepsilon=2\times 10^{-5}$. We define the total variance $w(k,\tau)=\sigma^{2}(k,\tau)\,\tau$ with forward $F_{\tau}$ and log-moneyness $k=\ln(K/F_{\tau})$. The total variance should be non-decreasing in $\tau$ at fixed $k$ if the IVS is calendar arbitrage-free. Tables 45 report violation rates, and Figure 5 visualizes butterfly spreads and total variance. Across Tables 45, SABR, SSVI, cubic spline, GP, and SABR-MTGP rarely exhibit butterfly violations on the evaluation grids; the few exceptions are small ($\leq 3.33\%$) and arise mainly for GP at long-term maturity and for cubic spline at near-term maturity. The NN baseline shows higher violation rates in several regimes, especially at long-term maturities. Figure 5 corroborates these findings: butterfly spreads of the proposed method are nonnegative, and total variance panels exhibit the expected increase in $w$ with maturity at fixed $k$ consistent with the absence of calendar arbitrage.
Table 4: Near-term butterfly violation rate ($\tau=0.3$). $\times$ denotes no violation.
Setting SABR SSVI Cubic Spline GP SABR-MTGP NN
Base $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Moderate Mean-Rev $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Low Mean-Rev $\times$ $\times$ 1.11% $\times$ $\times$ $\times$
High Vol-Regime $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Low Vol-Regime $\times$ $\times$ $\times$ $\times$ $\times$ 4.44%
Moderate Correlation $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Strong Correlation $\times$ $\times$ 2.22% $\times$ $\times$ 1.11%
Term Structure Up $\times$ $\times$ $\times$ $\times$ $\times$ 13.33%
Term Structure Down $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Mixed Regime $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Table 5: Long-term butterfly violation rate ($\tau=2.2$). $\times$ denotes no violation.
Setting SABR SSVI Cubic Spline GP SABR-MTGP NN
Base $\times$ $\times$ $\times$ 3.33% $\times$ 25.56%
Moderate Mean-Rev $\times$ $\times$ $\times$ $\times$ $\times$ 26.67%
Low Mean-Rev $\times$ $\times$ $\times$ 1.11% $\times$ 22.22%
High Vol-Regime $\times$ $\times$ $\times$ $\times$ $\times$ 37.78%
Low Vol-Regime $\times$ $\times$ $\times$ $\times$ $\times$ 14.44%
Moderate Correlation $\times$ $\times$ $\times$ $\times$ $\times$ 33.33%
Strong Correlation $\times$ $\times$ $\times$ $\times$ $\times$ 24.44%
Term Structure Up $\times$ $\times$ $\times$ $\times$ $\times$ 33.33%
Term Structure Down $\times$ $\times$ $\times$ 1.11% $\times$ 21.11%
Mixed Regime $\times$ $\times$ $\times$ $\times$ $\times$ 35.56%
(a) Butterfly spreads across strikes for three maturities (Base).
(b) Total variance $w(k,\tau)$ across methods (six panels, Base).
Figure 5: No-arbitrage test. Panel (a) visualizes butterfly spreads; most methods remain above the zero line within a small tolerance band, while NN and GP occasionally dip negative at mid/long maturities. Panel (b) shows six panels of total variance $w(k,\tau)$ curves for SABR, SSVI, Cubic Spline, GP, SABR-MTGP, and NN; for most methods the level of $w$ increases with maturity at fixed $k$ and the $k$-profiles are smooth, with minor departures mainly for $N$ and GP.
@@ 5.3 Learned Task Relationship
The MTGP framework learns the cross-task relationship between the SABR source and the Heston target. Figure 6 illustrates the learned task correlation, the task covariance, and the variance decomposition (task-specific proportion $\kappa^{2}_{\mathcal{Z}}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$ versus shared proportion $\sigma^{2}_{h}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$) of two representative settings.
(a) Term Structure Down Setting.
(b) Mixed Regime setting.
Figure 6: Learned task relationship for two representative Heston settings: (a) Term Structure Down and (b) Mixed Regime. Each panel shows the task correlation, covariance, and variance decomposition.
In the Term Structure Down case, the cross-task correlation is higher ($\approx 0.71$), indicating stronger transfer when the regime aligns with SABR’s structure. In the Mixed Regime, the lower correlation ($\approx 0.40$) implies greater reliance on task-specific variation. This adaptive behavior limits negative transfer while preserving the benefits of information transfer from SABR.
5.4 Sensitivity Analyses
5.4.1 Noise in the source task
We study how noise $\epsilon_{\mathcal{S}}\sim\mathcal{N}(0,\sigma_{\text{syn}}^{2})$ in the SABR-generated source task affects model accuracy. We inject zero-mean Gaussian noise with standard deviation $\sigma_{\text{syn}}\in\{0,\,0.005,\,0.01,\,0.015,\,0.02\}$ into the synthetic source implied volatilities, and evaluate overall accuracy on the same out-of-sample grid used elsewhere in this section. Figure 7 reports the Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) aggregated over the evaluation grid for three methods: SABR-MTGP, single-task GP, and a neural network pretrained on the source and finetuned on the target (NN).
Empirically, SABR-MTGP exhibits a clear U-shaped response: small amounts of noise reduce error and the best performance is attained around $0.01$, after which errors slowly increase. This behavior is consistent with the role of mild perturbations as a regularizer that reduces over-reliance on the synthetic dataset, while excessive noise reduces useful structural information. The NN baseline degrades monotonically with increasing $\sigma_{\text{syn}}$, reflecting its higher sensitivity to the quality of the pretraining dataset. Practically, these results suggest that modest source-task noise (here, $\sigma_{\text{syn}}$ in the range [0.005, 0.015][0.005,\,0.015]) can improve the robustness and generalization of SABR-MTGP.
Figure 7: Sensitivity of overall accuracy to Gaussian noise injected into the SABR source labels. Left: RMSE vs. noise level $\sigma_{\text{syn}}$. Right: MAE vs. $\sigma_{\text{syn}}$. SABR-MTGP benefits from modest noise (minimum near $\sigma_{\text{syn}}\!\approx\!0.01$), and NN degrades as $\sigma_{\text{syn}}$ increases.
5.4.2 Backbone parameter $\beta$
In the previous sections, we treat the SABR backbone $\beta$ as a hyperparameter for generating synthetic data. In this section, we verify that our conclusions do not hinge on its value. We repeat the Base experiment for $\beta\in\{0,\,0.25,\,0.5,\,0.75,\,1.0\}$, and compute out-of-sample errors on the same evaluation grid. Figure 8 shows that SABR-MTGP is insensitive to $\beta$ within a broad range, and outperforms GP in all cases. The neural network baseline is less accurate.
Figure 8: Effect of the SABR backbone parameter $\beta$ on overall accuracy in the Base setting. Left: RMSE vs. $\beta$. Right: MAE vs. $\beta$.
@@ 5.5 Summary
Experiments on Heston data show that SABR-MTGP effectively constructs the IVS. By combining SABR’s structural information with market observations in a multitask framework, it achieves better accuracy and stability than either GP or SABR alone, adaptively leveraging SABR insights according to their relevance to the target data.
@@ 6 Application to SPX Market Data
To assess the practical applicability of SABR-MTGP, we evaluate it on SPX option data from the OptionMetrics database dated August 1, 2023. This case study complements the controlled experiments using the Heston model by assessing the method’s behavior under actual market conditions.
@@ 6.1 Data
The dataset includes European call options on the SPX index. We preprocess the data by filtering for valid quotes, computing time-to-maturity, and computing implied volatilities from mid-prices. We use the zero-coupon yield curve for discounting and estimate the dividend yield via put-call parity. This gives us several hundred valid market implied volatility observations across different strikes and maturities up to about one year. Key statistics summarizing this dataset are in Table 6. This processed market dataset serves as the target task ($\mathcal{D}_{\mathcal{T}}$) for all methods.
@@ 6.2 Results
Figure 9 shows typical implied volatility slices for near-term ($\tau=0.2190$ years) and mid-term ($\tau=0.8898$ years) maturities.
@@
Figure 9: Implied volatility slices for SPX options on August 1, 2023. Comparison of SABR-MTGP, GP, and SABR against market data points. Residuals (Model - Market) are shown below each slice.
In both slices (Figure 9), all three methods generally capture the observed smile/skew pattern of the market. However, the residual plots reveal notable differences in fit quality. For the near-term slice (Figure 9(a)), SABR has systematic errors shown by the consistent pattern in its residuals. Both the GP and SABR-MTGP fit the observed market points better, especially around the at-the-money region. However, there is a notable difference in the deep in-the-money region (strikes 3250-4000). Although downward-sloping, the standard GP curve exhibits local irregularities that render it less financially plausible than the SABR-implied curve. The SABR-MTGP curve maintains a smoother, decreasing volatility in this data-sparse region. This produces a more financially realistic interpolation, likely because of the structural guidance from the SABR source task. For the mid-term slice (Figure 9(b)), where the smile is flatter and the market data is denser, all models achieve a very good fit. Notably, the SABR model fit improves significantly compared to the near-term slice. The residuals for both GP and SABR-MTGP are particularly small (generally within +/- 0.001), demonstrating excellent agreement with market observations.
@@
(d) SSVI
(e) Cubic Spline
(f) NN
Figure 10: SPX implied volatility surfaces constructed using six methods: GP, SABR, SABR-MTGP, SSVI, cubic spline, and a neural network. Black dots denote market observations.
Across the six surfaces in Figure 10, SABR-MTGP produces a smooth, well-behaved surface that tracks market levels across maturities and remains stable in sparse regions. The single-task GP is competitive overall but exhibits mild local undulations near the wings and toward the long end. The SABR and the neural network are smooth but can be too rigid around the wings. SSVI delivers a shape-constrained surface, whereas the cubic spline is generally smooth but shows occasional edge bending and a localized bulge at mid-maturities and ATM. Overall, the figures illustrate the intended trade-off: SABR-MTGP balances structural guidance and flexibility, preserving smoothness and plausibility while remaining close to the market points.
Figure 11: Butterfly spreads across strikes.
The butterfly spreads shown in Figure 11 remain nonnegative for SABR-MTGP, SABR, SSVI and cubic spline on the inspected slices, while GP and NN exhibit occasional dips near mid or long maturities. These diagnostics indicate that the surface generated by the proposed method is consistent with static no-arbitrage on the SPX sample.
@@ 7 Conclusion
In this paper, we proposed SABR-MTGP, an approach bridging structural and data-driven methods for IVS construction. The key idea is framing IVS construction as a multitask learning problem where we used structural information from the SABR model to improve predictions from sparse market data.
Our approach offered several contributions. First, we show how to combine financial models with machine learning techniques. We used SABR as a complementary information source within our multitask framework. Second, we developed a hierarchical Bayesian regularization for task embeddings that adaptively determined the level of information transfer. Third, our evaluation using Heston ground truth data showed this approach balanced structure and flexibility. It provided reliable predictions across market conditions and outperformed the standard Gaussian process, particularly in data-sparse regions. Improvements were competitive for longer maturities and extreme strikes, where market data was limited. Finally, an application to real SPX market data further illustrated the model’s ability to produce a smooth, reliable IVS that respects market observations while benefiting from structural guidance.
Our work has practical implications for quantitative finance. SABR-MTGP provides a more reliable tool for IVS construction, crucial for risk management. Hedging and OTC valuation require market-consistent dynamic models (e.g., local or stochastic local volatility) and Greek validation. These topics are outside the scope of this paper. By incorporating structural knowledge and empirical observations, it reduced model misspecification risk while maintaining adaptability to market realities. Future research could extend the multitask framework to other structural models or multiple models simultaneously.